OpenStudy (imactuallychris):

Find the difference. 4ab2 − 3ab2 1 a ab2 ab

OpenStudy (2500161):

4ab^2-3ab^2=ab^2(4-3)=ab^2

OpenStudy (imactuallychris):

thx!

OpenStudy (2500161):

Your welcome @imactuallychris